OpenStudy (anonymous):

Determine the limit. Correction 2.

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} {{4 - 7x} \over {2 + 3x}} = \lim_{x \rightarrow -\infty} {{4/x - 7} \over {2/x + 3}} = \lim_{x \rightarrow -\infty} {-7 \over 3}\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

The last one doesn't need the limit before it though :( (-7/3)

OpenStudy (anonymous):

Hahhaha! Sorry again.